what is the value of x to make 3x-5=2x+14
Answer by pmesler(52)

When you solve for x, you want to constantly simplify until all you're left with is x.
3x-5=2x+14.
To simplify this we need to get rid of the 5 on the left-hand side. We do this buy adding 5 from both sides. We add because we already have a negative 5 on the left side. By adding a positive 5 to a negative 5, the two cancel each other and we are left with zero. (Remember: what you do on one side of the equation you MUST do on the other side otherwise it wouldn't be an equation.)
3x-5(+5) = 2x + 14(+5)
3x + 0 = 2x + 19
3x = 2x + 19.

Now, we have one more step and we will find x. To get x alone we need to subtract 2x from each side. By doing this, the right side 2x's will cancel to zero and we will be left with 19 on the ride side and on the left side we will have the expression 3x-2x = 1x = x
So,
3x - 2x = 2x -2x + 19
x = 0 + 19
x = 19
Now let's check and see if this is correct by substituting x into the original equation.
3(19)-5 = 2(19)+14
52 = 52
Since both sides of the equation are equal x = 19 is the correct answer.